OpenStudy (anonymous):

mark works strictly on commission of his gross sales from sellling two different products. Last month his gross sales were $94,000. If he earns 4% commission on product A and 3% on product B what were his gross sales for each product if he earned $3480 in commission?

OpenStudy (anonymous):

system of equations first for the sales amount A + B = 94000 second for his commission .04(A) + .03(B) = 3480 just solve the system

OpenStudy (anonymous):

Sorry, I still don't get it. Do I set up the equation: 0.04 (94000) +0.03(94000-A) = 3489?

OpenStudy (anonymous):

close.... with a system of equations you want to solve one equation for one variable.. so lets solve the first one for B... we would get B = 94000-A now we go to the second one and plug in what we found for B so we would get .04(A) + .03(94000-A) = 3489 now we have an equation that we can solve since B is no longer involved... there is only A

OpenStudy (anonymous):

okay - got it! thank you! :-)